Fe2O3 + 3CO  --> 2Fe +  3CO2, how many grams of iron are produced when 7.47 grams of iron(III) oxide are reacted?


1
Expert's answer
2022-03-04T04:04:05-0500

Solution:

Calculate the moles of iron(III) oxide (Fe2O3):

The molar mass of Fe2O3 is 159.69 g/mol

Therefore,

Moles of Fe2O3 = (7.47 g Fe2O3) × (1 mol Fe2O3 / 159.69 g Fe2O3) = 0.0468 mol Fe2O3


Balanced chemical equation:

Fe2O3 + 3CO → 2Fe + 3CO2

According to stoichiometry:

1 mol of Fe2O3 produces 2 mol of Fe

Thus, 0.0468 mol of Fe2O3 produces:

(0.0468 mol Fe2O3) × (2 mol Fe / 1 mol Fe2O3) = 0.0936 mol Fe


Calculate the mass of iron (Fe):

The molar mass of Fe is 55.845 g/mol

Therefore,

Mass of Fe = (0.0936 mol Fe) × (55.845 g Fe / 1 mol Fe) = 5.227 g Fe = 5.23 g Fe

Mass of Fe = 5.23 g


Answer: 5.23 grams of iron (Fe) are produced